LASER-INDUCED FLUORESCENCE IN KLi

\underline{A. J. ROSS}, F. MARTIN, P. CROZET and M. AUBERT-FRÉCON, Laboratoire de Spectrométrie Ionique et Moléculaire (UMR 5579 CNRS), Bâtiment A. Kastler, Université Lyon I, Domaine Scientifique de la Doua, 69622 Villeurbanne Cedex, France; P. KOWALCZYK, Institute of Experimental Physics, Warsaw University , Ul. Hoża 69, 00-681 Warsaw, Poland; W. JASTRZȨBSKI, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warsaw, Poland.; and A. PASHOV, Universität Hannover, Institut für Quantenoptik, Welfengarten 1, 30167 Hannover, Germany.

Laser-induced fluorescence spectra of the B → X system of the potassium-lithium molecule have been recorded on a Fourier transform interferometer. The molecules were excited with a tuneable, single-mode cw dye laser operating with Rhodamine 110 or Coumarine 6 dye, and the spectra recorded in the range 12500 - 18500 cm^{−1} at a resolution of 0.05 cm^{−1}. Vibrational levels up to v = 46 are observed for the main isotopomer, ^{39}K^{7}Li. Ground state energies have been fitted to a Dunham polynomial expansion, and also directly to a pointwise potential curve. Both approaches fit the data with a root-mean-square error of 0.005 cm^{−1}. The data define the ground state potential curve accurately out to about 7 Å. The dissociation energy of the ground state D_{e} is estimated to be D_{e} = 6216 ± 2 cm^{−1}, and the highest vibrational level v_{max} = 50 for the main isotopomer. Very weak transitions around 12600 cm^{−1} are tentatively assigned as bands of the 2 ^{3}Σ^{+} → 1 ^{3}Σ^{+} transition, observed as collisionally induced fluorescence.